OpenStudy (anonymous):

find the values of a,b,c in the following identity: (a) a+b(x+2) +c(x+2)(x+3)=x^2+4x-2

OpenStudy (zarkon):

evaluate at -2 then -3

OpenStudy (sriram):

c=1 b=-1 a=-6

OpenStudy (sriram):

just compare coefficients of x^2 x and the constants

OpenStudy (lalaly):

\[a+bx+2b+c(x^2+5x+6) = x^2 +4x-2\]\[cx^2+(b+5)x + (a+2b+6c)=x^2 +4x-2\]\[c=1\]\[b+5=4 => b=-1\]\[a+2b+6c=-2 => a-2+6=-2 => a+4=-2 => a=-6\]

OpenStudy (anonymous):

thanks guys:)